OpenStudy (anonymous):

a riverboat travels 60 kilometers downstream in 6 hours, the return trip takes 10 hours/ find the speed of the boat in still water?

OpenStudy (anonymous):

current=c boat speed=a 6(a+c)=60 10(a-c)=60

OpenStudy (anonymous):

6a+6c=60 10a-10c=60

OpenStudy (anonymous):

Two variable two equation ;solve

OpenStudy (anonymous):

I got 8 and 2 x=8 y=2

OpenStudy (anonymous):

That's what I got

OpenStudy (anonymous):

so what speed is it 8 kilo or 2 kilo

OpenStudy (anonymous):

boat speed=8 water current=2

OpenStudy (anonymous):

I get it now thanks